OpenStudy (anonymous):

I know how can x^2+xy+y^2=0 be solved but how can I solve 3x^2+8xy-3y^2=2012 ? there is 2012 instead of 0

OpenStudy (raden):

there are 2 equations ?

OpenStudy (anonymous):

no, only 3x^2+8xy-3y^2=2012

OpenStudy (raden):

with x, y were integer numbers right ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, first make factoring it will giving us (3x-y)(x+3y) = 2012 right ?

OpenStudy (anonymous):

yes. But how did you realize that? :)

OpenStudy (raden):

it's simple, try and errors :p

OpenStudy (raden):

i meant, with try rearranged to get the factor of them

OpenStudy (anonymous):

ok ok. so then?

OpenStudy (raden):

u must find the factors of 2012 too :)

OpenStudy (anonymous):

2012=2*3*337

OpenStudy (raden):

2012 cant divided by 3 :)

OpenStudy (anonymous):

ah, I was doing it on my mind :D and I confused 1006 with 1011...

OpenStudy (anonymous):

ok, can you continue and solve it? :)

OpenStudy (raden):

i can, but i must let u doing together :p

OpenStudy (raden):

remember, the rule's from "Openstudy" :p

OpenStudy (anonymous):

I'm new here :D

OpenStudy (raden):

but, im not sure u cant find the factors of 2012, right ? :)

OpenStudy (anonymous):

2^2*503

OpenStudy (raden):

yes, nice.. :)

OpenStudy (raden):

but it should be 2012 = +- {1,2,4,503,1006,2012} right ?

OpenStudy (anonymous):

yes, I just found prime factors :D

OpenStudy (raden):

ok, now any posibles that can we doing from there, for an example : (3x-y)(x+3y) = 1 * 2012 , it means 3x-y=1 and x+3y=2012, with elimination and subtitution method, we can find the solution of them, it getting us x=201.5 and y=... but, exactly x is not an integer number, so not satisfies from the 1st posible equation, can u find the others ?

OpenStudy (anonymous):

ok, I got it. thanks ;)

OpenStudy (raden):

ok, glad to hear u can understand this problem.. :) very welcome

OpenStudy (raden):

but, as compare, u can look the answer in wolfram : http://www.wolframalpha.com/input/?i=3x^2%2B8xy-3y^2%3D2012